The thermite reaction is very exothermic, the heat of the reaction sufficient to melt the iron that is produced by the reaction:

2Al + Fe2O3 Al2O3 + 2Fe H=?

a) calculate the standard enthalpy change for this reaction?
b) if 40.0 g of iron (III) oxide we’re reacted with excess aluminum, how much heat would be produced ?
c) of the heat produced was used to heat 2.00L of water at 23.0C, what would the final temperature of the water be?
1
Expert's answer
2019-07-17T06:17:18-0400

a) deltaH(Al2O3)=-1675.5 kJ/mol,

deltaH(Fe2O3)=-824.2 kJ/mol,

deltaH(reaction)=deltaH(Al2O3)-deltaH(Fe2O3)=-1675.5+824.2=-851.3kJ/mol


b)n(Fe2O3)=m/M=40/159.7=0.25 mol,

deltaH=n*deltaH(reaction)=0.25*851.3=212.8 kJ


c) Q=c(water)*m(water)*(t2-t1), t2= t1+(Q/c(water)*m(water))= 48.4 C